OpenStudy (anonymous):

The cost C (in dollars) of manufacturing x chairs at Martina's Furniture Factory is given by the function c(x)=0.4x^2-160x+34,108. What is the minimum cost of manufacturing chairs?

OpenStudy (er.mohd.amir):

d/dx{c(x)}=0.4*2*x-160 0.8x-160=0 x=160/0.8=200

OpenStudy (er.mohd.amir):

200 is min since d^2/dx^2=0.8 is +ve

OpenStudy (anonymous):

18,108 Is what they wanted but thanks

OpenStudy (er.mohd.amir):

put x=200 in c(x)=0.4*200*200160*200+34108=18108

OpenStudy (er.mohd.amir):

that is at x=200 c(x) is min and on any other it give more value of 18108